OpenStudy (anonymous):

I have no idea how to do this \[ \sqrt{98w^{12}x^{68}}\] Simplify the expression assuming x and w may be negative

hero (hero):

I would rather assume that w and x is positive

OpenStudy (anonymous):

\[\sqrt{98}\sqrt{w^{12} x^{68}}=7 \sqrt{2} w^6 x^{34} \]